# A Probabilistic Framework for LLM Hallucination Detection via Belief Tree Propagation

### v1

@@ 5 Conclusion
6 Acknowledgements
The work of Bairu Hou and Shiyu Chang was partially supported by National Science Foundation (NSF) Grant IIS-2207052 and NSF Grant IIS-2302730. The computing resources used in this work were partially supported by the Accelerate Foundation Models Research program of Microsoft.
7 Broader Impact
@@ A.1 Implementation Details
We preprocess the data in Wikibio-GPT3 [4] and FELM-Science [15] before evaluating our method and baselines. The data preprocessing includes two steps. First, we perform decontextualization since the sentences in a response generated by the LLM might be context-dependent and contain pronouns or noun phrases that cannot be understood without full context. Take the first data instance from Wikibio-GPT3 for demonstration, which is a biography of “John Russell Reynolds” generated by an LLM. The second sentence within the biography requiring hallucination detection is “He was born in London, the son of a barrister, and was educated at Eton College and Trinity College, Cambridge.” Without the context, we cannot identify the truthfulness of this sentence due to the pronoun. Therefore, we first decontextualize the sentences in the two datasets by prompting gpt-3.5-turbo-0125. The prompt is available in Figure 9.
@@
We use the vLLM [44] to perform the inference of Llama-3-8B-Instruct. We use a single NVIDIA A100 80GB PCIe GPU to evaluate the performance and report the time cost in Figure 4. For the emission probability estimation (Table 1), we use the first 120 examples in the Wikibio-GPT3 dataset (50% of it) to compute the empirical estimation of the emission probability. The model we use is gpt-3.5-turbo-0125. Then we transfer the estimated emission probabilities to the remaining examples in Wikibio-GPT3 and other datasets for performance evaluation. For Llama-3-8B-Instruction, we also use the same estimated emission probabilities. Additionally, since we set the model confidence on child nodes generated by statement correction as 1.0 rather than probe the model’s confidence, we maintain an individual emission probability for those child nodes ($p(S_{u}=1|Z_{u}=1)=0.8$ and $p(S_{u}=1|Z_{u}=0)=0.2$). During the belief tree construction process, we add several constraints to boost the efficiency. Specifically, we only apply statement decomposition to the root node, assuming the child nodes generated by the LLM is atomic statements that only contain one aspect of information. Also, we do not expand nodes generated by statement correction, as they are homogeneous to their parent node.
@@
We follow the default configuration of each baselines. For SelfCheckgpt, we sample 20 additional responses for hallucination detection. Similarly, we also sample 20 different answers for chain-of-thought prompting and aggregate them using self-consistency [45]. For Maieutic-Prompting, we use their prompts for the CREAK datasets to generate belief trees in our evaluation.
A.2 Examples of Belief Trees
@@
We visualize several constructed belief trees as well as how our method leverage the inconsistencies in model’s beliefs for hallucination detection. We start with a simple example from the Wikibio-GPT3 dataset [4] shown in Figure 5. Starting from the root node about Adiele Afigbo, the first layer contains 3 child nodes generated by statement decomposition. However, despite the statement is wrong, the model (gpt-3.5-turbo-0125) assign a high confidence score to both the original statement and the child nodes decomposed from the root node. Then, our method further generate child nodes for node 1,2, and 3. We display the child nodes of node 3 in the figure, which is generated by statement correction. With the three different statements about the date of death of that person, we decrease the model’s confidence on node 3 and finally correct the model’s belief on the root node. Note that in statement correction, we directly set the confidence of the generated child nodes as 1.0. For this example, if we query the model’s confidence scores on node 4, 5, and 6, we will get confidence scores 0.97, 0.91, 0.88, respectively, which will still contradicts to their parent node.
@@
Another example of the belief tree is shown in Figure 6, which mainly consists of child nodes generated by supportive and contradictory premises. At the root node, the model assigns a high confidence score to the statement about freezing point of water from the FELM dataset [15]. After the statement decomposition, the inconsistency is triggered due to model’s low confidence on node 2. Furthermore, if we continuously generate premises for node 1, we get two additional child nodes (node 3 and 4) that are contradictory to their parent node. However, the model still assigns a high confidence to node 3 and 4. Within the belief propagation framework, the conditional probability of node 1 being true given the observations on node 3 and 4 will be decreased. Similarly, this effect will propagate to the root node and lead to a low posterior probability of the root node being true.
@@
We then show one failure case of our method in Figure 7. The target statements comes from the FactCheckGPT dataset [43]. Our method first generation two supportive premises with high confidence. These two child nodes are indeed reasonable and can support the opinion that “Linux operating system has a widespread adoption.”, thus contradicting to the root statement and further decreasing the model’s confidence on the root node. We hypothesize that the original statement, “Linux adoption has limited compared to other operating systems like Windows and macOS” lacks sufficient specificity and could be interpreted from multiple perspectives. For example, it could refer to market share in personal computing versus cloud computing domains, making the ground-truth label nondeterministic.
A.3 Inferring the Posterior Probability of $Z_{0}$
The posterior probability $p(Z_{0}=1|\{S_{u}\})$ can be computed by:
$p(Z_{0}=1|\{S_{u}\})=\frac{p(Z_{0}=1,\{S_{u}\})}{p(Z_{0}=1,\{S_{u}\})+p(Z_{0}=0,\{S_{u}\})},$ (2)
@@ A.3 Inferring the Posterior Probability of $Z_{0}$
Therefore, the key of the inference on the belief tree is to compute the two joint probabilities $p(Z_{0}=1,\{S_{u}\})$ and $p(S_{1}=\texttt{False},\bar{\bm{X}}_{1}=\bar{\bm{x}}_{1})$. Following the conditional independence assumption in the hidden Markov model, the truthfulness (i.e,, states of hidden variables) of each node is determined by its parent node in the tree and the transition probability. Also, the model’s confidence (i.e., states of observed variables) on each node is determined by the state of the corresponding variable and the emission probability. With such an assumption, they can be decomposed as:
$p(Z_{0}=1,\{S_{u}\})$ $=p(\{S_{u}\}|Z_{0}=1)*p(Z_{0}=1),$ (3)
where $p(Z_{0}=1)$ is the prior probability of the statement being true (note that we set it to be 0.5 in this paper). $p(\{S_{u}\}|Z_{0}=1)$ is the conditional probability of all observed variables (exclude the root node) given the root node being true. Recall that we introduce a notation $\beta(z,u)=p(S_{\mathcal{T}(u)})|Z_{u}=z$ to represent such conditional probabilities, where $S_{{\mathbb{T}}(u)}$ refers to the confidence scores (observed variables) of all nodes on the subtree rooted at node $u$. Therefore, the conditional probability $p(\{S_{u}\}|Z_{0}=1)$ can be denoted as $\beta(1,0)$ (node 00, hidden variable $Z_{0}=1$).
Without loss of generality, we discuss below how to compute $\beta(z,u)$ for an arbitrary node $u$ in the tree. The computation for $\beta(1,0)$ and $\beta(0,0)$ can be performed in exactly the same way. Specifically, we can further decompose $\beta(z,u)$ as:
$p(S_{\mathcal{T}(u)})|Z_{u}=z)$ =$=\left\{\prod_{v\in\bm{{\mathcal{C}}}(u)}p(S_{\mathcal{T}(v)}|Z_{u}=z)\right\}p(S_{u}|Z_{u}=z).$ (4)
The probability $p(S_{\mathcal{T}(v)}|Z_{u}=z)$ can be decomposed as:
$\centering p(S_{\mathcal{T}(v)}|Z_{u}=z)$ $=\sum_{k\in\{0,1\}}p(S_{\mathcal{T}(v)}|Z_{v}=k)*p(Z_{v}=k|Z_{u}=z)$ (5)
$=\sum_{k\in\{0,1\}}\beta(v,k)*p(Z_{v}=k|Z_{u}=z).$
Therefore, the conditional probability $\beta(z,u)=p(S_{\mathcal{T}(u)})|Z_{u}=z)$ can be computed in a recursive manner:
$\beta(z,u)=\left\{\prod_{v\in\bm{{\mathcal{C}}}(u)}\sum_{k\in\{0,1\}}\beta(v,k)*\underbrace{p(Z_{v}=k|Z_{u}=z)}_{\text{Transition Probability}}\right\}\underbrace{p(S_{u}|Z_{u}=z)}_{\text{Emission Probability}}.$ (6)
There are two types of probabilities in the above equation. First, $p(S_{u}|Z_{u}=z)$ is the “emission probability" at node $u$. Second, $p(Z_{v}=k|Z_{u}=z)$ is the “transition probability" from node $u$ to its child node $v$. The recursive computation in Equation 6 will starts from the leaf nodes. When $v$ is the leaf node in the tree, $\beta(v,k)$ is actually $p(S_{v}|Z_{v}=k)$, which is the emission probability at node $v$. Such a process propagates from bottom to up. Finally, we can compute $\beta(1,0)$ as well as the joint probability $p(Z_{0}=1|\{S_{u}\})$ according to Equation 3.
@@
The above computation of the posterior probability is based on the assumption of conditional independence. Assume the node $u$ have two child nodes, $v_{1}$ and $v_{2}$. Then we consider the transition probabilities from $u$ to $v_{1}$ and from $u$ to $v_{2}$ independently. However, this does not hold for the child nodes generated by statement decomposition, in which the truthfulness of the child nodes are influenced by their parent node simultaneously. If $u$ is true, then we know that both two child nodes are also true. In contrast, if $u$ is false, we can only infer that at least one of the child nodes are false. To handle such a case, we revise the probability computation in Equation 4 accordingly. Assume node $u$ has $m$ child nodes. The probability $\beta(z,u)=p(S_{\mathcal{T}(u)})|Z_{u}=z)$ can be computed as:
$\centering p(S_{\mathcal{T}(u)})|Z_{u}=z)$ $p(S_{u}|Z_{u}=z)\cdot\sum_{Z_{\mathcal{C}(u)}\in\{0,1\}^{m}}p(Z_{\mathcal{C}(u)}|Z_{u}=z)\prod_{v\in\mathcal{C}(u)}\beta(Z_{v},v),$ (7)
@@ A.4 Summary of the Algorithm
2:Output: The posterior probability $p(Z_{0}=1|S_{\mathcal{T}(0)})$.
@@
29:$p(Z_{0}=1|S_{\mathcal{T}(0)})=\beta(1,0)/(\beta(1,0)+\beta(0,0))$
We summarize the pipeline of our method in Algorithm 1. The root node of the belief tree is the given statement $u_{0}$. During the belief tree construction process, we maintain a set $\mathcal{N}$ that contains all current leaf nodes of the belief tree. We recursively expand each leaf node $u\in\mathcal{N}$ by its child nodes $\mathcal{C}(u)$ and add these new child nodes to the belief tree if they are logically-connected to the current leaf node. These generated child nodes then become new leaf nodes. They will also be added to $\mathcal{N}$ and be expanded until the maximum depth is reached. Given the constructed belief tree and the emission and transition probabilities estimated from a held-out dataset, the posterior probability $\beta(Z_{0}=1|S_{\mathcal{T}(0)})$ can be then computed in a recursive manner, which will be further used to predict the truthfulness of the given statement.
A.5 License of the Datasets Used
FactCheckGPT is under Apache-2.0 license. FELM is under CC-BY-NC-SA-4.0 license, and Wikibio-GPT3 is under CC-BY-SA-3.0 license. We have cited them accordingly in the main paper.
A.6 Prompt

### v2

@@ 5 Conclusion
6 Societal Impact and limitations
@@
Acknowledgments
The work of Bairu Hou and Shiyu Chang was partially supported by National Science Foundation (NSF) Grant IIS-2338252, NSF Grant IIS-2207052, and NSF Grant IIS-2302730. The work of Jacob Andreas is supported by a Sloan Fellowship and the NSF under grant IIS-2238240. The computing resources used in this work were partially supported by the Accelerate Foundation Models Research program of Microsoft and CAIS Compute Cluster of Center for AI Safety.
@@ A.1 Implementation Details
We preprocess the data in Wikibio-GPT3 Manakul et al. (2023) and FELM-Science Chen et al. (2023b) before evaluating our method and baselines. The data preprocessing includes two steps. First, we perform decontextualization since the sentences in a response generated by the LLM might be context-dependent and contain pronouns or noun phrases that cannot be understood without full context. Take the first data instance from Wikibio-GPT3 for demonstration, which is a biography of “John Russell Reynolds” generated by an LLM. The second sentence within the biography requiring hallucination detection is “He was born in London, the son of a barrister, and was educated at Eton College and Trinity College, Cambridge.” Without the context, we cannot identify the truthfulness of this sentence due to the pronoun. Therefore, we first decontextualize the sentences in the two datasets by prompting gpt-3.5-turbo-0125. The prompt is available in Figure 9.
@@
We use the vLLM Kwon et al. (2023) to perform the inference of Llama-3-8B-Instruct. We use a single NVIDIA A100 80GB PCIe GPU to evaluate the performance and report the time cost in Figure 4. For the emission probability estimation (Table 1), we use the first 120 examples in the Wikibio-GPT3 dataset (50% of it) to compute the empirical estimation of the emission probability. The model we use is gpt-3.5-turbo-0125. Then we transfer the estimated emission probabilities to the remaining examples in Wikibio-GPT3 and other datasets for performance evaluation. For Llama-3-8B-Instruction, we also use the same estimated emission probabilities.
For the emission probability, we use first 120 examples of the Wikibio-GPT3 dataset as the validation set and estimate the emission probability on it using gpt-3.5-turbo-0125. The estimated emission probabilities are in Table 1 and will used on the other testing examples of Wikibio-GPT3 and other two datasets. Additionally, since we set the model confidence on child nodes generated by statement correction as 1.0 rather than probe the model’s confidence, we find setting an individual set of emission probabilities for those statements improves the performance. Therefore, we tune the emission probabilities for child nodes generated by statement correction on the validation set. The emission probabilities for those child nodes we use are $p(S_{u}=1\mid Z_{u}=1)=0.8$ and $p(S_{u}=1\mid Z_{u}=0)=0.2$. During the belief tree construction process, we add several constraints to boost the efficiency. Specifically, we only apply statement decomposition to the root node, assuming the child nodes generated by the LLM is atomic statements that only contain one aspect of information. Also, we do not expand nodes generated by statement correction, as they are homogeneous to their parent node.
@@
We follow the default configuration of each baselines. For SelfCheckGPT, we sample 20 additional responses for hallucination detection. Similarly, we also sample 20 different answers for chain-of-thought prompting and aggregate them using self-consistency Wang et al. (2022). For Maieutic-Prompting, we use their prompts for the CREAK datasets to generate belief trees in our evaluation. For all methods, we use the scikit-learn package to compute the evaluation metrics including AUROC, AUC-PR, and F1 score.
A.2 Additional Results
AUROC AUC-PR F1 Acc
Prior Confidence 67.7 63.2 71.6 68.1
MSP 70.5 73.3 69.3 57.4
Perplexity 66.0 63.4 70.0 59.3
CCP 75.8 78.1 72.4 69.5
BTProp 76.5 75.5 73.3 70.0
Since the datasets in the above experiments consist of text pre-generated by another model for hallucination detection, we also include texts generated by the hallucination detection model itself to further evaluate the effectiveness of our method. Specifically, we first generate biographical data using the prompt from Wikibio-GPT3 dataset with Llama-3-8b-instruct, following the “claim-level uncertainty quantification” setup of LM-Polygraph (Fadeeva et al., 2023). Then we use the same model to detect the hallucinations in its generations, where the ground-truth label for each claim is determined by FactScore (Min et al., 2023). We evaluate the performance of our method and additional baselines including CCP (Fadeeva et al., 2024), maximum sequence probability (MSP), and Perplexity. The ground-truth label for each claim is determined by FactScore. The results are summarized in Table 5.
Similar to the observation in our main experiments, our method performs competitively against these state-of-the-art uncertainty quantification techniques. Specifically, our approach achieves the highest AUROC, F1 scores, and accuracy, demonstrating its ability to reliably detect hallucinations. Please note that our method is not originally designed for the “claim-level” hallucination detection, where the biography data are decomposed into atomic claims for fact-checking. Nevertheless, our method still achieves competitive performance.
A.3 Examples of Belief Trees
@@
We visualize several constructed belief trees as well as how our method leverage the inconsistencies in model’s beliefs for hallucination detection. We start with a simple example from the Wikibio-GPT3 dataset Manakul et al. (2023) shown in Figure 5. Starting from the root node about Adiele Afigbo, the first layer contains 3 child nodes generated by statement decomposition. However, despite the statement is wrong, the model (gpt-3.5-turbo-0125) assign a high confidence score to both the original statement and the child nodes decomposed from the root node. Then, our method further generate child nodes for node 1,2, and 3. We display the child nodes of node 3 in the figure, which is generated by statement correction. With the three different statements about the date of death of that person, we decrease the model’s confidence on node 3 and finally correct the model’s belief on the root node. Note that in statement correction, we directly set the confidence of the generated child nodes as 1.0. For this example, if we query the model’s confidence scores on node 4, 5, and 6, we will get confidence scores 0.97, 0.91, 0.88, respectively, which will still contradicts to their parent node.
@@
Another example of the belief tree is shown in Figure 6, which mainly consists of child nodes generated by supportive and contradictory premises. At the root node, the model assigns a high confidence score to the statement about freezing point of water from the FELM dataset Chen et al. (2023b). After the statement decomposition, the inconsistency is triggered due to model’s low confidence on node 2. Furthermore, if we continuously generate premises for node 1, we get two additional child nodes (node 3 and 4) that are contradictory to their parent node. However, the model still assigns a high confidence to node 3 and 4. Within the belief propagation framework, the conditional probability of node 1 being true given the observations on node 3 and 4 will be decreased. Similarly, this effect will propagate to the root node and lead to a low posterior probability of the root node being true.
@@
We then show one failure case of our method in Figure 7. The target statements comes from the FactCheckGPT dataset Wang et al. (2023). Our method first generation two supportive premises with high confidence. These two child nodes are indeed reasonable and can support the opinion that “Linux operating system has a widespread adoption.”, thus contradicting to the root statement and further decreasing the model’s confidence on the root node. We hypothesize that the original statement, “Linux adoption has limited compared to other operating systems like Windows and macOS” lacks sufficient specificity and could be interpreted from multiple perspectives. For example, it could refer to market share in personal computing versus cloud computing domains, making the ground-truth label nondeterministic.
A.4 Inferring the Posterior Probability of $Z_{0}$
The posterior probability $p(Z_{0}=T\mid\{S_{u}\})$ can be computed by:
$\frac{p(Z_{0}=T,\{S_{u}\})}{p(Z_{0}=T,\{S_{u}\})+p(Z_{0}=F,\{S_{u}\})},$
@@ A.4 Inferring the Posterior Probability of $Z_{0}$
Therefore, the key of the inference on the belief tree is to compute the two joint probabilities $p(Z_{0}=T,\{S_{u}\})$ and $p(S_{1}=F,\bar{\bm{X}}_{1}=\bar{\bm{x}}_{1})$. Following the conditional independence assumption in the hidden Markov model, the truthfulness (i.e,, states of hidden variables) of each node is determined by its parent node in the tree and the transition probability. Also, the model’s confidence (i.e., states of observed variables) on each node is determined by the state of the corresponding variable and the emission probability. With such an assumption, they can be decomposed as:
$p(Z_{0}=T,\{S_{u}\})$ $=p(\{S_{u}\}\mid Z_{0}=T)*p(Z_{0}=T),$ (2)
where $p(Z_{0}=T)$ is the prior probability of the statement being true (note that we set it to be 0.5 in this paper). $p(\{S_{u}\}\mid Z_{0}=T)$ is the conditional probability of all observed variables (exclude the root node) given the root node being true. Recall that we introduce a notation $\beta(z,u)=p(S_{\mathcal{T}(u)}|Z_{u}=z)$ to represent such conditional probabilities, where $S_{\mathcal{T}(u)}$ refers to the confidence scores (observed variables) of all nodes on the subtree rooted at node $u$. Therefore, the conditional probability $p(\{S_{u}\}\mid Z_{0}=T)$ can be denoted as $\beta(T,0)$ (node 00, hidden variable $Z_{0}=T$).
Without loss of generality, we discuss below how to compute $\beta(z,u)$ for an arbitrary node $u$ in the tree. The computation for $\beta(T,0)$ and $\beta(F,0)$ can be performed in exactly the same way. Specifically, we can further decompose $\beta(z,u)$ as:
$p(S_{\mathcal{T}(u)})\mid Z_{u}=z)$ $=p(S_{u}\mid Z_{u}=z)\cdot$ (3)
$\left\{\prod_{v\in\bm{{\mathcal{C}}}(u)}p(S_{\mathcal{T}(v)}\mid Z_{u}=z)\right\}.$
The probability $p(S_{\mathcal{T}(v)}\mid Z_{u}=z)$ can be decomposed as:
$p(S_{\mathcal{T}(v)}\mid Z_{u}=z)$ (4)
$=\sum_{k\in\{T,F\}}p(S_{\mathcal{T}(v)}\mid Z_{v}=k)*p(Z_{v}=k\mid Z_{u}=z)$
$=\sum_{k\in\{T,F\}}\beta(v,k)*p(Z_{v}=k\mid Z_{u}=z).$
Therefore, the conditional probability $\beta(z,u)=p(S_{\mathcal{T}(u)})\mid Z_{u}=z)$ can be computed in a recursive manner:
$\beta(z,u)=\underbrace{p(S_{u}\mid Z_{u}=z)}_{\text{Emission Probability}}\cdot$ (5)
$\left\{\prod_{v\in\bm{{\mathcal{C}}}(u)}\sum_{k\in\{T,F\}}\beta(v,k)*\underbrace{p(Z_{v}=k\mid Z_{u}=z)}_{\text{Transition Probability}}\right\}.$
There are two types of probabilities in the above equation. First, $p(S_{u}\mid Z_{u}=z)$ is the “emission probability" at node $u$. Second, $p(Z_{v}=k\mid Z_{u}=z)$ is the “transition probability" from node $u$ to its child node $v$. The recursive computation in Equation 5 will starts from the leaf nodes. When $v$ is the leaf node in the tree, $\beta(v,k)$ is actually $p(S_{v}\mid Z_{v}=k)$, which is the emission probability at node $v$. Such a process propagates from bottom to up. Finally, we can compute $\beta(t,0)$ as well as the joint probability $p(Z_{0}=T\mid\{S_{u}\})$ according to Equation 2.
@@
The above computation of the posterior probability is based on the assumption of conditional independence. Assume the node $u$ have two child nodes, $v_{1}$ and $v_{2}$. Then we consider the transition probabilities from $u$ to $v_{1}$ and from $u$ to $v_{2}$ independently. However, this does not hold for the child nodes generated by statement decomposition, in which the truthfulness of the child nodes are influenced by their parent node simultaneously. If $u$ is true, then we know that both two child nodes are also true. In contrast, if $u$ is false, we can only infer that at least one of the child nodes are false. To handle such a case, we revise the probability computation in Equation 3 accordingly. Assume node $u$ has $m$ child nodes. The probability $\beta(z,u)=p(S_{\mathcal{T}(u)})\mid Z_{u}=z)$ can be computed as:
$p(S_{\mathcal{T}(u)})\mid Z_{u}=z)=p(S_{u}\mid Z_{u}=z)\tilde{\beta}(u,z),\text{where}$
$\tilde{\beta}(u,z)=\sum_{Z_{\mathcal{C}(u)}\in\{T,F\}^{m}}p(Z_{\mathcal{C}(u)}\mid Z_{u}=z)\prod_{v\in\mathcal{C}(u)}\beta(Z_{v},v)$
@@ A.5 Summary of the Algorithm
2:Output: The posterior probability $p(Z_{0}=1\mid S_{\mathcal{T}(0)})$.
@@
29:$p(Z_{0}=1\mid S_{\mathcal{T}(0)})=\beta(1,0)/(\beta(1,0)+\beta(0,0))$
We summarize the pipeline of our method in Algorithm 1. The root node of the belief tree is the given statement $u_{0}$. During the belief tree construction process, we maintain a set $\mathcal{N}$ that contains all current leaf nodes of the belief tree. We recursively expand each leaf node $u\in\mathcal{N}$ by its child nodes $\mathcal{C}(u)$ and add these new child nodes to the belief tree if they are logically-connected to the current leaf node. These generated child nodes then become new leaf nodes. They will also be added to $\mathcal{N}$ and be expanded until the maximum depth is reached. Given the constructed belief tree and the emission and transition probabilities estimated from a held-out dataset, the posterior probability $\beta(Z_{0}=1\mid S_{\mathcal{T}(0)})$ can be then computed in a recursive manner, which will be further used to predict the truthfulness of the given statement.
A.6 Submission Checklist
FactCheckGPT is under Apache-2.0 license. FELM is under CC-BY-NC-SA-4.0 license, and Wikibio-GPT3 is under CC-BY-SA-3.0 license. We have cited them accordingly in the main paper. All of these datasets are in English. These datasets are created to evaluate the hallucination detection performance, and our usages are consistent with their intended use. We also manually check the datasets and confirm that they do not contains any information that names or uniquely identifies individual people or offensive content.
We leverage ChatGPT to help develop some of the prompt we use in the experiments and improve the writing.
A.7 Prompt